BBO Discussion Forums: All you have to do is take 9 tricks - BBO Discussion Forums

Jump to content

Page 1 of 1
  • You cannot start a new topic
  • You cannot reply to this topic

All you have to do is take 9 tricks

#1 User is offline   Siegmund 

  • Alchemist
  • PipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 1,764
  • Joined: 2004-June-15
  • Gender:Male
  • Location:Beside a little lake in northwestern Montana
  • Interests:Creator of the 'grbbridge' LaTeX typesetting package.

Posted 2013-January-27, 00:42

♠AK43
♥AQJ2
♦J3
♣763

♠986
♥943
♦AKQ98
♣K4

You don't have to like your auction, but here it is:

South deals, favorable, and opens a 12-14 1NT.
West overcalls 2C promising the black suits.
Partner propels you into 3NT after using Stayman.

The opening lead is a club to the jack and king. You are obviously set unless you win nine in a row.

Are you going to play for the drop in diamonds, or will you play a diamond to the jack and finesse on the way back, expecting West to be short after his vulnerable 2-suited overcall?
0

#2 User is offline   Stephen Tu 

  • PipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 4,155
  • Joined: 2003-May-14

Posted 2013-January-27, 01:43

The drop has to be a huge favorite. If West has only 9 cds in blacks like most people promise then drop is still best play in diamond suit. If West promise 10 cds in blacks, finesse in diamond become favorite for 5 trick in that suit, but not by a ton, much less than your chance of finding west with Kx hearts.
1

#3 User is offline   EricK 

  • PipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 2,303
  • Joined: 2003-February-14
  • Location:England

Posted 2013-January-27, 02:54

I have 8 on top don't I (4♦, 1♣, 2♠, 1♥)? Playing for the drop in ♦ gives me the chance to try the ♥ finesse if the ♦ don't break. The combined odds of that seem larger than that of finding East with ♦T. The exact odds are hard to determine - although the vulnerable overcall makes it more likely that West has short ♦, it might also make it more likely he has ♥K (depending somewhat on how constructive the opps are over a weak NT).
0

#4 User is offline   gnasher 

  • Andy Bowles
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,993
  • Joined: 2007-May-03
  • Gender:Male
  • Location:London, UK

Posted 2013-January-27, 03:44

View PostStephen Tu, on 2013-January-27, 01:43, said:

The drop has to be a huge favorite. If West has only 9 cds in blacks like most people promise then drop is still best play in diamond suit. If West promise 10 cds in blacks, finesse in diamond become favorite for 5 trick in that suit, but not by a ton, much less than your chance of finding west with Kx hearts.


It wouldn't have to be Kx. Any holding containing ♥K would be enough, because West would be squeezed when you cashed the top diamonds. For example, with 4315 he'd have to discard a spade and two clubs, but you'd take a heart finesse, cash ♠AK to confirm the count, then give up a spade.
... that would still not be conclusive proof, before someone wants to explain that to me as well as if I was a 5 year-old. - gwnn
0

#5 User is offline   FrancesHinden 

  • Limit bidder
  • PipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 8,483
  • Joined: 2004-November-02
  • Gender:Female
  • Location:England
  • Interests:Bridge, classical music, skiing... but I spend more time earning a living than doing any of those

Posted 2013-January-27, 04:17

A good rule in this sort of position is that the finesse is always worse odds that you think it is (if you see what I mean)
So my intuition is to play for the drop, even without the extra chance of the HK onside.

p.s. you don't need any subtle squeeze if the HK is onside, you've got 4 diamonds, 1 club, 2 hearts, 2 spades
0

#6 User is offline   rhm 

  • PipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 3,092
  • Joined: 2005-June-27

Posted 2013-January-27, 04:17

View PostSiegmund, on 2013-January-27, 00:42, said:

The opening lead is a club to the jack and king. You are obviously set unless you win nine in a row.

Appearances are deceptive
Return a club.
Of course West may switch immediately to a heart, but he also may not.
As far as I can see as long as West does not have six clubs all your chances remain intact even if West switches to hearts.

Rainer Herrmann
0

#7 User is offline   FrancesHinden 

  • Limit bidder
  • PipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 8,483
  • Joined: 2004-November-02
  • Gender:Female
  • Location:England
  • Interests:Bridge, classical music, skiing... but I spend more time earning a living than doing any of those

Posted 2013-January-27, 04:20

Now I've calculated it, I make the drop a 29:23 favourite if West has 9 black suit cards
0

#8 User is offline   Antrax 

  • PipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 2,458
  • Joined: 2011-March-15
  • Gender:Male

Posted 2013-January-27, 04:34

rhm, can you explain the merit of playing a club? I'm thinking even my partner and I would manage to cash four clubs and find a good switch in that situation, since E will have plenty of time to signal.
0

#9 User is offline   gnasher 

  • Andy Bowles
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,993
  • Joined: 2007-May-03
  • Gender:Male
  • Location:London, UK

Posted 2013-January-27, 04:58

View PostAntrax, on 2013-January-27, 04:34, said:

rhm, can you explain the merit of playing a club? I'm thinking even my partner and I would manage to cash four clubs and find a good switch in that situation, since E will have plenty of time to signal.

Suppose that the defence cashed the clubs and played a heart. Declarer could keep all his original chances by playing the ace, then diamonds from the top. If diamonds didn't come in but West had ♥K, he would be squeezed in the majors.

If the defence switched to a heart at trick three, again declarer could play ♥A followed by the top diamonds. If diamonds didn't come in but West had ♥K, he would be three-suit squeezed.

Where Rainer's idea gains is if the defence cashes the clubs but you decide to play for RHO to have ♥K. After a heart switch, win the ace and then cash the top spades to squeeze East.
... that would still not be conclusive proof, before someone wants to explain that to me as well as if I was a 5 year-old. - gwnn
0

#10 User is offline   rhm 

  • PipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 3,092
  • Joined: 2005-June-27

Posted 2013-January-27, 06:52

View Postgnasher, on 2013-January-27, 04:58, said:

View PostAntrax, on 2013-January-27, 04:34, said:

rhm, can you explain the merit of playing a club? I'm thinking even my partner and I would manage to cash four clubs and find a good switch in that situation, since E will have plenty of time to signal.

Suppose that the defence cashed the clubs and played a heart. Declarer could keep all his original chances by playing the ace, then diamonds from the top. If diamonds didn't come in but West had ♥K, he would be squeezed in the majors.

If the defence switched to a heart at trick three, again declarer could play ♥A followed by the top diamonds. If diamonds didn't come in but West had ♥K, he would be three-suit squeezed.

Where Rainer's idea gains is if the defence cashes the clubs but you decide to play for RHO to have ♥K. After a heart switch, win the ace and then cash the top spades to squeeze East.

To add to this:

If clubs are 5-3 West does not know that East has a third club and that he should switch to hearts at trick three, whoever has the heart king.
He does not know that your source of tricks is diamonds not hearts and from his point of view it looks certainly more likely from bidding and play that declarer has a third club, where the club return would be even more indicated.
If West cashes 2 or 3 club tricks before switching you can safely take the heart finesse.
If West cashes 4 clubs and does not switch to hearts you should play for a show-up double squeeze with hearts as the common threat.
Technically this is not a double squeeze since only one can have the heart king and control the heart suit, but you do not care.
Whoever has the king will be squeezed when you cash the spades and run the diamonds.

Playing a club back just gives the defense a chance to go wrong and in my experience ( I have not so much experience with worldclass defense) they often will.

Rainer Herrmann
1

#11 User is offline   cherdano 

  • 5555
  • PipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 9,520
  • Joined: 2003-September-04
  • Gender:Male

Posted 2013-January-27, 06:57

Does this really offset the chances of West having 6 clubs?
The easiest way to count losers is to line up the people who talk about loser count, and count them. -Kieran Dyke
0

#12 User is offline   rhm 

  • PipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 3,092
  • Joined: 2005-June-27

Posted 2013-January-27, 07:28

View Postcherdano, on 2013-January-27, 06:57, said:

Does this really offset the chances of West having 6 clubs?

How likely is it that West has six or more clubs when he has spades as well?
If you look at all distribution which would qualify from 5-4 in the black suits onwards this will happen less than one occurrence in five.
Of course that West had enough to interfere and his lead changes these odds a bit.
However, even more important, in this case given that West has at least six clubs and a spade suit, the chance that diamonds will be 1-5 and the heart king offside increases substantially.

I believe that Bridge is not played double dummy and it is very worthwhile to look at a problem from a defender's point of view.
I am prepared to invest something to induce an error even if alternative lines are not hopeless.

Rainer Herrmann
0

#13 User is offline   gnasher 

  • Andy Bowles
  • PipPipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 11,993
  • Joined: 2007-May-03
  • Gender:Male
  • Location:London, UK

Posted 2013-January-27, 07:55

View Postcherdano, on 2013-January-27, 06:57, said:

Does this really offset the chances of West having 6 clubs?

For it to matter, he'd have to have ♥K or ♦10 as well as being 4-6. If we assume that he'd have bid on any hand with the right shape, given what we know about the high cards, I think the chance of his being 4-6 with one of these cards is about 5%.

There are 5 ways that clubs can be AQ10xxx-Jx or A10xxxx-QJ, and 10 ways that they can be AQ10xx-Jxx or A10xxx-QJx . After dealing all the black cards except for one club, the vacant spaces are 4:9. Hence the probability of any possible 6-2 break is 4/13 * 5/15 = 4/39.

When he's 4-6, the vacant spaces are 3:9, so the chance that he has either ♥K or ♦10 is 3/12 + (9/12) * 3/11 ~= 45%. That figure multipled by 4/39 is about 5%

Obviously the assumption in my first paragraph may not be valid. And my calculations may not be right - I've already edited them three times (originally I made it 11%, then 10, then 7, and now 5).

This post has been edited by gnasher: 2013-January-27, 09:07

... that would still not be conclusive proof, before someone wants to explain that to me as well as if I was a 5 year-old. - gwnn
0

#14 User is offline   Siegmund 

  • Alchemist
  • PipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 1,764
  • Joined: 2004-June-15
  • Gender:Male
  • Location:Beside a little lake in northwestern Montana
  • Interests:Creator of the 'grbbridge' LaTeX typesetting package.

Posted 2013-January-27, 11:53

I am enjoying the discussion. I got to watch declarer guess wrong, but could tell it was a reasonably close decision, and couldn't tell what I would have done if I had been playing it. I am glad others enjoyed considering the problem too.

The thought of a club immediately crossed my mind too but I thought there was enough chance of a 6-2 club break I didn't mention it in the OP.
0

#15 User is offline   nige1 

  • 5-level belongs to me
  • PipPipPipPipPipPipPipPipPip
  • Group: Advanced Members
  • Posts: 9,128
  • Joined: 2004-August-30
  • Gender:Male
  • Location:Glasgow Scotland
  • Interests:Poems Computers

Posted 2013-January-27, 20:04

View PostSiegmund, on 2013-January-27, 00:42, said:

North ♠ AK43 ♥ AQJ2 ♦ J3 ♣ 763
South ♠ 986 ♥ 943 ♦ AKQ98 ♣ K4
You don't have to like your auction, but here it is: South deals, favorable, and opens a 12-14 1NT. West overcalls 2C promising the black suits. Partner propels you into 3NT after using Stayman. The opening lead is a club to the jack and king. You are obviously set unless you win nine in a row. Are you going to play for the drop in diamonds, or will you play a diamond to the jack and finesse on the way back, expecting West to be short after his vulnerable 2-suited overcall?

View Postrhm, on 2013-January-27, 06:52, said:

If clubs are 5-3 West does not know that East has a third club and that he should switch to hearts at trick three, whoever has the heart king.
As usual, Rainer's line of exiting in ♣ is excellent. Against Rainer, however, when LHO is thrown in at trick two, unless he can cash five ♣ tricks, he should exit immediately with a ♥ -- on general spoil-sport principles -- i.e What Rainer seems to want us to do is likely to be bad for us . Perhaps, you should just win ♣K, and cash as many ♦s as you can. Unless LHO has precisely a small singleton ♦ you are home. And even if ♦s are worth only four tricks, you still retain some chances.
0

Page 1 of 1
  • You cannot start a new topic
  • You cannot reply to this topic

2 User(s) are reading this topic
0 members, 2 guests, 0 anonymous users